OpenStudy (anonymous):

A bus traveling 40 mi/h left Freetown at noon. A car following the bus at 60 mi/h left Freetown at 1:30 P.M. At what time did the car catch up with the bus?

Parth (parthkohli):

OK - so if the bus travelled for \(x\) hours, then the car travelled for \(x - 1.5\) hours.

Parth (parthkohli):

\[\rm distance = speed \times time \]

Parth (parthkohli):

\[40\times x = 60 \times (x - 1.5)\]